# Differentially Private Zeroth-Order Methods for Scalable Large Language Model Fine-tuning

### v4

@@ Abstract.
We provide theoretical analysis for both proposed methods. We conduct extensive empirical analysis on both encoder-only masked language model and decoder-only autoregressive language model, achieving impressive results in terms of scalability and utility (compared with DPZero (Zhang et al., 2023), DP-ZOPO improves $4.5\%$ on SST-5, $5.5\%$ on MNLI with RoBERTa-Large and 9.2% on CB, 3.9% on BoolQ with OPT-2.7B when $\epsilon=4$).
1. Introduction
@@
Pretrained Large Language Models (LLMs), scaling up to unprecedented sizes, have demonstrated remarkable potential in their capabilities of understanding and processing natural languages with human-like proficiency (Radford et al., 2018) (Radford et al., 2019). This has prompted a rapid surge in demand to harness the power of pretrained LLMs, particularly open-sourced series like OPT (Zhang et al., 2022), llama (Touvron et al., 2023) and GPT (Radford et al., 2019), to boost performance across a wide range of downstream tasks. Fine-tuning is one of the most fundamental and dominant approaches for adapting pretrained LLMs to specific downstream tasks, which has been proven effective in (Ziegler et al., 2019) (Sun et al., 2019). Fine-tuning starts with the publicly accessible checkpoint of the selected model and continues to train the model for several epochs based on the task-specific dataset (referred to as the fine-tuning dataset hereafter). While appearing a simple process at first glance, the sheer scale of nowadays LLMs introduces significant scalability issues for fine-tuning, e.g., incurring prohibitive memory footprint, which complicates the design of training methods even in nonprivate fine-tuning (Zhang et al., 2023).
@@
It is widely recognized that a finetuned LLM can leak sensitive information (Carlini et al., 2019) (Carlini et al., 2021) from its fine-tuning dataset, which is considered private and valuable in certain downstream application areas related to finance, healthcare (Carlini et al., 2021). Therefore, private fine-tuning of pretrained LLMs has become a pressing need due to the escalating privacy concerns associated with the growing popularity of LLMs (Yu et al., 2021a). Differential privacy (DP) (Dwork, 2006) is a widely adopted mathematical framework that ensures a rigorous privacy protection guarantee by introducing calibrated random perturbations. A long-standing research theme in DP is the pursuit of an ideal tradeoff between privacy and utility, which arises because the random perturbation for privacy-preserving purpose will inevitably degrade the utility. DP LLM fine-tuning faces even intensified tension as it has to handle the tradeoff among privacy, utility, and scalability.
@@
Most of the existing DP LLM fine-tuning methods (Li et al., 2021) (He et al., 2022) build upon the seminal work of Differentially Private SGD (DP-SGD) (Abadi et al., 2016), which clips the gradient vector per training sample, injects random Gaussian noise after clipping, and tracks the privacy budget across iterations by the moments accountant technique. These DP-SGD-based methods mitigate the scalability issue in roughly two lines of effort. The first line (Liu et al., 2021) (Bu et al., 2022) (He et al., 2022) aims to reduce the computational cost of the per-sample clipping of the gradient vector that is known to incur heavy computational and memory burden. For instance, (Liu et al., 2021) introduce group clipping to replace the per-sample clip, (Bun et al., 2014) propose book keep (BK) strategy. The second line (Feng et al., 2023) incorporates DP-SGD with parameter-efficient fine-tuning techniques in LLMs, which limits the number of trainable parameters. The intuition is that the utility-privacy tradeoff degrades quickly with respect to the number of trainable parameters since the DP perturbations need to be injected into all dimensions of the gradient vector. For instance, (Yu et al., 2021a) (Li et al., 2021) consider LoRA, adapters, and compacter, which either train two smaller factors to approximate the LLM parameter update or insert additional trainable modules while freezing the pretrained LLM parameters.
@@
Although the DP-SGD-based methods mentioned above have made the best effort to squeeze performance out of DP-SGD, the inefficiency inherent in SGD continues to hinder achieving a satisfactory “privacy-utility-scalability” tradeoff for LLM fine-tuning. That is, the gradient calculation requires caching all intermediate activations during the forward pass and gradients during the backward pass, which leads to prohibitive memory usage for LLM fine-tuning, i.e., consuming up to 12×12$\times$12 × the memory usage required for inference (Malladi et al., 2023). Driven by such limitation, recent nonprivate LLM fine-tuning methods turn to zeroth-order methods, which circumvent the scalability issues associated with gradient computations by approximating them using two inferences (Malladi et al., 2023) (Zhao et al., 2024) (Zhang et al., 2024). In nonprivate LLM fine-tuning, the effectiveness of zeroth-order methods has been validated through both theoretical and empirical studies, which suggests a promising direction for developing the DP LLM fine-tuning method.
@@
Our contributions. First, we focus on the synergy between the DP random perturbations and the gradient approximation error of zeroth-order estimation to the true gradient, across different training stages. Specifically, we propose to dynamically adjust the ZO scale parameter that controls the gradient approximation error and divide the DP LLM fine-tuning into stages. We propose DP-ZOSO, the first stagewise algorithm for differentially private zeroth-order fine-tuning to our best knowledge. In earlier stages, the gradient approximation error is controlled to be smaller, and together with a larger learning rate, it allows the LLM fine-tuning to approach the optimum more quickly. In the later stages, we need to deliberately increase the gradient approximation error, together with a decreased learning rate, offering a stabilization effect on the fine-tuning trajectory. Our theoretical analysis demonstrates that this stagewise DP zeroth-order fine-tuning strategy provides an improved convergence rate.
@@
Third, we conduct extensive empirical evaluations to corroborate the superiority of our proposed DP zeroth-order methods for LLM fine-tuning. We utilize four different open-source LLMs, including masked language model (RoBERTa-large) and autoregressive language model (OPT-2.7B) for downstream tasks such as sentiment classification, natural language inference, and topic classification (including datasets like SST-2, SST-5, SNLI, MNLI, and TREC). Our method exhibits strong scalability and performs well across all models and datasets. Our method has enhanced the model’s performance greatly (improves $4.5\%$ on SST-5, $5.5\%$ on MNLI with RoBERTa-Large and 9.2% on CB, 3.9% on BoolQ with OPT-2.7B when $\epsilon=4$) compared with existing differentially private zeroth-order fine-tuning method.
@@
  • We extensively experiment with our methods, providing empirical evidence of their scalability and utility. DP-ZOPO outperforms parameter-efficient method (DP prefix tuning) by 10.9% on MNLI and 7.6% on TREC when $\epsilon=4$ with RoBERTa-large. DP-ZOPO also outperforms the zeroth-order method (DPZero) by 5.5% on MNLI and 4.4% on TREC when $\epsilon=4$ with RoBERTa-large (see Figure 1).
@@
Figure 1. Experiments on RoBERTAa-large. We report zero-shot, DPZero(Zhang et al., 2023), DP-ZOPO, DP-ZOSO and fine-tuning (FT) with full parameter and prefix-tuning. DP-ZOSO and DP-ZOSO both outperform zero-shot, FT (prefix) with much less memory. DP-ZOPO far outperforms DPZero, and DP-ZOSO and approaches FT (Detailed numbers in Table 2).
2. Background and Related Work
2.1. Private Parameter Efficient Fine Tuning
@@
As LLMs scale up, full-parameter fine-tuning becomes impractical due to the requirement for extensive GPU memory. Things only get worse with privacy, which leads to overheads in terms of running time, memory usage, and most importantly, accuracy. The magnitude of noise introduced to a model due to DP grows as the model size increases, which can result in poor performance of LLMs. Parameter-efficient fine-tuning methods reduce memory consumption by updating just a fraction of the network parameters. Differentially private prompt tuning (Li et al., 2023) (Duan et al., 2023) has emerged as a simple and parameter-efficient solution for steering LLMs to downstream tasks. DP fine-tuning with Reparametrized Gradient Perturbation (Yu et al., 2021b) (Li et al., 2021) computes the low-dimension projected gradient without computing the gradient itself, significantly reducing computation cost. Besides parameter-efficient fine-tuning, there also exist other memory-efficient methods.
2.2. DP Zeroth-Order Fine-Tuning
@@
Zeroth-order fine-tuning methods decrease memory consumption by replacing backpropagation with two inferences (Malladi et al., 2023) (Zhang et al., 2024), each done with the same random perturbation with flipped signs. Recently, there have been several research that explores differential private zeroth-order fine-tuning: a workshop paper (Zhang et al., 2023) and a “work in progress” paper on arXiv (Tang et al., 2024). We stress that our work differs from them in three key aspects: 1) Regarding algorithm design, all these works merely consider constant scheduling of hyperparameters and have not explored parameter sparsification to further boost scalability like us; 2) Regarding theoretical analysis: (Tang et al., 2024) did not provide any theoretical analysis for the utility. (Zhang et al., 2023)’s utility analysis did not consider the dynamic scheduling of the hyperparameters that render the analysis more involved; 3) Regarding empirical analysis: (Zhang et al., 2023) did not provide any empirical studies. While (Tang et al., 2024) also considered LLM finetuing, our empirical study is more comprehensive by considering both RoBERTa-large and OPT models, along with classification and multiple tasks. Moreover, targeting the sparsity inherent in LLMs, we propose DP-ZO fine-tuning with pruning and achieve better performance both theoretically and empirically.
2.3. DP-SGD with Sparsification
@@
Due to the low-rankness or the sparsity of neural networks, pruning has been widely used as a dimensionality reduction method to improve the scalability of DP-SGD (Adamczewski and Park, 2023) (Luo et al., 2021) (Mireshghallah et al., 2022). Adamczewski et. al. (Adamczewski et al., 2023) proposed parameter freezing pruning method: pre-prune the network and update those selected using DP-SGD. They (Adamczewski and Park, 2023) further propose parameter selection: select which parameters to update at each step of training and update only those selected using DP-SGD. They use public data for freezing or selecting parameters to avoid privacy loss incurring in these steps. However, there has been no study about stagewise pruning method with dynamic pruning rates in differentially private zeroth-order fine-tuning.
3. Preliminary
3.1. LLM Fine-Tuning
@@
Definition 3.1 (LLM Fine-tuning).
@@
(1) $\min\limits_{\boldsymbol{\theta}\in\mathcal{R}^{d}}\{f(\boldsymbol{\theta},\mathcal{D})\}:=\frac{1}{n}\sum_{i=1}^{n}f(\boldsymbol{\theta},x_{i}),$
@@
Definition 3.2 (Stochastic gradient descent).
@@
(2) $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta\cdot\nabla f(\boldsymbol{\theta}_{t-1},\xi_{t}),$
@@
3.2. Differential Privacy
Differential privacy aims to provide a mathematical definition for the privacy guarantees of an algorithm or system.
Definition 3.3 (Differential Privacy (Dwork et al., 2006)).
@@ 3.2. Differential Privacy
(3) $\mathbb{P}\left[\mathcal{A}(X)\in s\right]\leq e^{\epsilon}*\mathbb{P}\left[\mathcal{A}(X^{\prime})\in s\right]+\delta,$
@@
Differentially Private Stochastic Gradient Descent (DPSGD) is a privacy-preserving variant of SGD. It is designed to train models while providing strong guarantees of differential privacy.
Definition 3.4 (DP-SGD (Abadi et al., 2016)).
DP-SGD is a differentially private optimizer with clipping threshold $C$, noise scale $\sigma$ and learning rate $\eta$ that updates parameters as
(4) $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta\cdot\frac{1}{m}\left(\sum_{i}\hat{g}(\boldsymbol{\theta}_{t-1},x_{i})+N(0,\sigma^{2}C^{2}\textbf{I}_{d})\right),$
@@
The following three lemmas are frequently used in the privacy analysis of DP-SGD.
Lemma 3.5 (Privacy amplification by subsampling (Kasiviswanathan et al., 2011)).
Let $\mathcal{A}:\mathcal{X}^{r_{2}}\rightarrow\mathcal{S}$ be a $(\epsilon,\delta)$-DP machanism. Then, the mechanism $\mathcal{A}^{\prime}:\mathcal{X}^{r_{1}}\rightarrow\mathcal{S}$ defined by $\mathcal{A}^{\prime}=\mathcal{A}\circ\operatorname{samp}_{r_{1},r_{2}}$ is $\left(\epsilon^{\prime},\delta^{\prime}\right)$-DP, where
(5) $\epsilon^{\prime}=\log\left(1+q\left(e^{\epsilon}-1\right)\right),\delta^{\prime}=q\delta,q=\frac{r_{2}}{r_{1}}.$
In particular, when $\epsilon<1,\mathcal{A}^{\prime}$ is $(\mathcal{O}(q\epsilon),q\delta)$-DP.
Lemma 3.6 (Strong composition (Dwork et al., 2014)).
Let $\mathcal{A}_{1},\dots,\mathcal{A}_{T}$ be $T$ adaptive ($\epsilon,\delta$)-DP mechanisms, where $\epsilon,\delta\geq 0$. Then, for any $\delta^{\prime}\geq 0$, the composed mechanism $\mathcal{A}=(\mathcal{A}_{1},...,\mathcal{A}_{T})$ is ($\hat{\epsilon},\hat{\delta}$)-DP, where
(6) $\hat{\epsilon}=\sqrt{2T\log\left(\frac{1}{\delta^{\prime}}\right)}\epsilon+T\epsilon\left(e^{\epsilon}-1\right),\hat{\delta}=T\delta+\delta^{\prime}.$
Moments accountant method tracks the accumulation of privacy loss over multiple iterations. By using the moments accountant, it is possible to estimate the overall privacy loss over a sequence of operations, which provides a tighter privacy analysis compared with strong composition Lemma 3.6.
Lemma 3.7 (Moments Accountant (Abadi et al., 2016)).
There exisi constant $c_{1}$ and $c_{2}$ so that given the sampling probability $q=m/n$ and the number of steps $T$, for any $\epsilon<c_{1}q^{2}T$, DP-SGD is $(\epsilon,\delta)$-differentially private for any $\delta>0$ if we choose
(7) $\sigma\geq c_{2}\frac{q\sqrt{T\log(1/\delta)}}{\epsilon}.$
3.3. Zeroth-Order Optimization

### v5

@@ Abstract
We provide theoretical analysis for both proposed methods. We conduct extensive empirical analysis on both encoder-only masked language model and decoder-only autoregressive language model, achieving impressive results in terms of scalability and utility regardless of the class of tasks (compared with DPZero [1], DP-ZOPO improves $4.5\%$ on SST-5, $5.5\%$ on MNLI with RoBERTa-Large and 9.2% on CB, 3.9% on BoolQ with OPT-2.7b when $\epsilon=4$, demonstrates more significant enhancement in performance on more complicated tasks).
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 24 - 28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.241214 www.ndss-symposium.org
I Introduction
@@
Pretrained Large Language Models (LLMs), scaling up to unprecedented sizes, have demonstrated remarkable potential in their capabilities of understanding and processing natural languages with human-like proficiency [2] [3]. This has prompted a rapid surge in demand to harness the power of pretrained LLMs, particularly open-sourced series like OPT [4], llama [5] and GPT [3], to boost performance across a wide range of downstream tasks. Fine-tuning is one of the most fundamental and dominant approaches for adapting pretrained LLMs to specific downstream tasks, which has been proven effective in [6] [7]. Fine-tuning starts with the publicly accessible checkpoint of the selected model and continues to train the model for several epochs based on the task-specific dataset (referred to as the fine-tuning dataset hereafter). While appearing a simple process at first glance, the sheer scale of nowadays LLMs introduces significant scalability issues for fine-tuning, e.g., incurring prohibitive memory footprint, which complicates the design of training methods even in nonprivate fine-tuning [8].
@@
It is widely recognized that a finetuned LLM can leak sensitive information [9] [10] from its fine-tuning dataset, which is considered private and valuable in certain downstream application areas related to finance, healthcare [10]. Therefore, private fine-tuning of pretrained LLMs has become a pressing need due to the escalating privacy concerns associated with the growing popularity of LLMs [11]. Differential privacy (DP) [12] is a widely adopted mathematical framework that ensures a rigorous privacy protection guarantee by introducing calibrated random perturbations. A long-standing research theme in DP is the pursuit of an ideal tradeoff between privacy and utility, which arises because the random perturbation for privacy-preserving purpose will inevitably degrade the utility. DP LLM fine-tuning faces even intensified tension as it has to handle the tradeoff among privacy, utility, and scalability.
@@
Most of the existing DP LLM fine-tuning methods [13] [14] build upon the seminal work of Differentially Private SGD (DP-SGD) [15], which clips the gradient vector per training sample, injects random Gaussian noise after clipping, and tracks the privacy budget across iterations by the moments accountant technique. These DP-SGD-based methods mitigate the scalability issue in roughly two lines of effort. The first line [16] [17] [14] aims to reduce the computational cost of the per-sample clipping of the gradient vector that is known to incur heavy computational and memory burden. For instance, [16] introduce group clipping to replace the per-sample clip, [18] propose book keep (BK) strategy. The second line [19] incorporates DP-SGD with parameter-efficient fine-tuning techniques in LLMs, which limits the number of trainable parameters. The intuition is that the utility-privacy tradeoff degrades quickly with respect to the number of trainable parameters since the DP perturbations need to be injected into all dimensions of the gradient vector. For instance, [11] [13] consider LoRA, adapters, and compacter, which either train two smaller factors to approximate the LLM parameter update or insert additional trainable modules while freezing the pretrained LLM parameters.
@@
Although the DP-SGD-based methods mentioned above have made the best effort to squeeze performance out of DP-SGD, the inefficiency inherent in SGD continues to hinder achieving a satisfactory “privacy-utility-scalability” tradeoff for LLM fine-tuning. That is, the gradient calculation requires caching all intermediate activations during the forward pass and gradients during the backward pass, which leads to prohibitive memory usage for LLM fine-tuning, i.e., consuming up to 12×12$\times$12 × the memory usage required for inference [20]. Driven by such limitation, recent nonprivate LLM fine-tuning methods turn to zeroth-order methods, which circumvent the scalability issues associated with gradient computations by approximating them using two inferences [20] [21] [22]. In nonprivate LLM fine-tuning, the effectiveness of zeroth-order methods has been validated through both theoretical and empirical studies, which suggests a promising direction for developing the DP LLM fine-tuning method.
@@
Our contributions. First, we focus on the synergy between the DP random perturbations and the gradient approximation error of zeroth-order estimation to the true gradient, across different training stages. Specifically, we propose to dynamically adjust the ZO scale parameter that controls the gradient approximation error and divide the DP LLM fine-tuning into stages. We propose DP-ZOSO, the first stagewise algorithm for DP-ZO fine-tuning to our best knowledge. In earlier stages, the gradient approximation error is controlled to be smaller, and together with a larger learning rate, it allows the LLM fine-tuning to approach the optimum more quickly. In the later stages, we need to deliberately increase the gradient approximation error, together with a decreased learning rate, offering a stabilization effect on the fine-tuning trajectory. Our theoretical analysis demonstrates that this stagewise DP-ZO fine-tuning strategy provides an improved convergence rate.
@@
Third, we conduct extensive empirical evaluations to corroborate the superiority of our proposed DP zeroth-order methods for LLM fine-tuning. We utilize four different open-source LLMs, including masked language model (RoBERTa-large) and autoregressive language model (OPT-1.3b, OPT-2.7b) for downstream tasks such as sentiment classification, natural language inference, topic classification (including datasets like SST-2, SST-5, SNLI, MNLI, and TREC), question answering (SQuAD), translation (Europarl) and summarization (XSum and CNN/DM). Our method exhibits strong scalability and performs well across all models and datasets. Our method has enhanced the model’s performance greatly (improves $4.5\%$ on SST-5, $5.5\%$ on MNLI with RoBERTa-Large and 9.2% on CB, 3.9% on BoolQ with OPT-2.7b when $\epsilon=4$) compared with existing differentially private zeroth-order fine-tuning method. Similar results can be observed in the tasks of question-answering, translation and summarization tasks.
Figure 1: Experiments on RoBERTAa-large. We report zero-shot, DPZero [1], DP-ZOPO, DP-ZOSO and DP full-parameter fine-tuning (FT) and DP prefix-tuning (FT-prefix). DP-ZOSO and DP-ZOSO both outperform zero-shot, FT-prefix with much less memory. DP-ZOPO far outperforms DPZero, and DP-ZOSO and approaches FT (Detailed numbers in Table II).
@@
  • We extensively experiment with our methods, providing empirical evidence of their scalability and utility. DP-ZOPO outperforms parameter-efficient method (DP prefix tuning) by 10.9% on MNLI and 7.6% on TREC when $\epsilon=4$ with RoBERTa-large. DP-ZOPO also outperforms DPZero by 5.5% on MNLI and 4.4% on TREC when $\epsilon=4$ with RoBERTa-large (see Figure 1). More results pertaining to a broader range of tasks are presented in Section V-H.
@@
II Background and Related Work
II-A Private Parameter Efficient Fine-Tuning
@@
As LLMs scale up, full-parameter fine-tuning becomes impractical due to the requirement for extensive GPU memory. Things only get worse with privacy, which leads to overheads in terms of running time, memory usage, and most importantly, accuracy. The magnitude of noise introduced to a model due to DP grows as the model size increases, which can result in poor performance of LLMs. Parameter-efficient fine-tuning methods reduce memory consumption by updating just a fraction of the network parameters. Differentially private prompt tuning [23] [24] has emerged as a simple and parameter-efficient solution for steering LLMs to downstream tasks. DP fine-tuning with Reparametrized Gradient Perturbation [25] [13] computes the low-dimension projected gradient without computing the gradient itself, significantly reducing computation cost. Besides parameter-efficient fine-tuning, there also exist other memory-efficient methods.
II-B DP Zeroth-Order Fine-Tuning
@@
Zeroth-order fine-tuning methods decrease memory consumption by replacing backpropagation with two inferences [20] [22], each done with the same random perturbation with flipped signs. Recently, there have been several research that explores differential private zeroth-order fine-tuning: a workshop paper [1] and a “work in progress” paper on arXiv [26]. We stress that our work differs from them in three key aspects: 1) Regarding algorithm design, all these works merely consider constant scheduling of hyperparameters and have not explored parameter sparsification to further boost scalability like us; 2) Regarding theoretical analysis: [26] did not provide any theoretical analysis for the utility. [1]’s utility analysis did not consider the dynamic scheduling of the hyperparameters that render the analysis more involved; 3) Regarding empirical analysis: [1] did not provide any empirical studies. While [26] also considered LLM finetuing, our empirical study is more comprehensive by considering both RoBERTa-large and OPT models, along with serval complicated tasks. Moreover, targeting the sparsity inherent in LLMs, we propose DP-ZO fine-tuning with pruning and achieve better performance both theoretically and empirically.
II-C DP-SGD with Sparsification
@@
Due to the low-rankness or the sparsity of neural networks, pruning has been widely used as a dimensionality reduction method to improve the scalability of DP-SGD [27] [28] [29]. Adamczewski et. al. [30] proposed parameter freezing pruning method: pre-prune the network and update those selected using DP-SGD. They [27] further propose parameter selection: select which parameters to update at each step of training and update only those selected using DP-SGD. They use public data for freezing or selecting parameters to avoid privacy loss incurring in these steps. However, there has been no study about stagewise pruning method with dynamic pruning rates in differentially private zeroth-order fine-tuning.
III Preliminary
III-A LLM Fine-Tuning
@@
Definition 1 (LLM Fine-tuning).
@@
$\min\limits_{\boldsymbol{\theta}\in\mathcal{R}^{d}}\{f(\boldsymbol{\theta},\mathcal{D})\}:=\frac{1}{n}\sum_{i=1}^{n}f(\boldsymbol{\theta},x_{i}),$ (1)
@@
Definition 2 (Stochastic gradient descent).
@@
$\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta\cdot\nabla f(\boldsymbol{\theta}_{t-1},\xi_{t}),$ (2)
@@
III-B Differential Privacy
Differential Privacy (DP) aims to provide a rigorous mathematical definition for privacy guarantees.
Definition 3 (Differential Privacy [31]).
@@ III-B Differential Privacy
$\mathbb{P}\left[\mathcal{A}(X)\in s\right]\leq e^{\epsilon}*\mathbb{P}\left[\mathcal{A}(X^{\prime})\in s\right]+\delta,$ (3)
@@
DP-SGD is a privacy-preserving variant of SGD, designed to train with strong guarantees of differential privacy.
Definition 4 (DP-SGD [15]).
DP-SGD optimizes with clipping threshold $C$, noise scale $\sigma$, learning rate $\eta$ and updates parameters as
$\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta\cdot\frac{1}{m}\left(\sum_{i}\hat{g}(\boldsymbol{\theta}_{t-1},x_{i})+N(0,\sigma^{2}C^{2}\textbf{I}_{d})\right),$ (4)
@@
Moments accountant method tracks the accumulation of privacy loss over multiple iterations. By using the moments accountant, it is possible to estimate the overall privacy loss over a sequence of operations.
Lemma 1 (Moments Accountant (Poisson Subsampling) [15]).
There exist constant $c_{1}$ and $c_{2}$ so that given the sampling probability $q=m/n$ and the number of steps $T$, for any $\epsilon<c_{1}q^{2}T$, DP-SGD is $(\epsilon,\delta)$-differentially private for any $\delta>0$ if we choose
$\sigma\geq c_{2}\frac{q\sqrt{T\log(1/\delta)}}{\epsilon}.$ (5)
Lemma 2 (Privacy amplification by Poisson Subsampling and Uniform Sampling Without Replacement [32]).
Due to the convenience of unifrom sampling in convergence analysis, uniform sampling is also adopted in [33] [34].
  • Poisson Subsampling. Let $\mathcal{A}:\mathcal{X}\rightarrow\mathcal{S}$ be a $(\epsilon,\delta)$-DP machanism. Then, the mechanism $\mathcal{A}^{\prime}:\mathcal{X}\rightarrow\mathcal{S}$ defined by $\mathcal{A}^{\prime}=\mathcal{A}\circ\mathrm{samp}_{\tau}^{\mathrm{po}}$ where the poisson subsampling mechanism $\mathrm{samp}_{\tau}^{\mathrm{po}}$ takes a set $x$ and outputs a sample $y$ from the distribution $w=\mathrm{samp}_{\tau}^{\mathrm{po}}(x)$ supported on all set $y\subseteq x$ given by $w(y)=\tau^{|y|}(1-\tau)^{|x|-|y|}$. The mechanism $\mathcal{A}^{\prime}$ is $\left(\epsilon^{\prime},\delta^{\prime}\right)$-DP, where
    $\epsilon^{\prime}=\log\left(1+\tau\left(e^{\epsilon}-1\right)\right),\delta^{\prime}=\tau\delta.$ (6)
  • Uniform Sampling without Replacement. Let $\mathcal{A}:\mathcal{X}^{r_{2}}\rightarrow\mathcal{S}$ be a $(\epsilon,\delta)$-DP machanism. Then, the mechanism $\mathcal{A}^{\prime}:\mathcal{X}^{r_{1}}\rightarrow\mathcal{S}$ defined by $\mathcal{A}^{\prime}=\mathcal{A}\circ\mathrm{samp}_{r_{2}/r_{1}}^{\mathrm{wo}}$ where the subsampling mechanism $\mathrm{samp}_{r_{2}/r_{1}}^{\mathrm{wo}}$ takes a set $x$ of size $r_{1}$ and outputs a sample from the uniform distribution $w=\mathrm{samp}_{r_{2}/r_{1}}^{\mathrm{wo}}(x)$ over all sebsets $y\subseteq x$ of size $r_{2}\leq r_{1}$. The mechanism $\mathcal{A}^{\prime}$ is $\left(\epsilon^{\prime},\delta^{\prime}\right)$-DP, where
    $\epsilon^{\prime}=\log\left(1+\frac{r_{2}}{r_{1}}\left(e^{\epsilon}-1\right)\right),\delta^{\prime}=\frac{r_{2}}{r_{1}}\delta.$ (7)
    Since Poisson subsampling and uniform sampling without replacement share the same order of privacy amplification. Following the privacy analysis of Moments Accountant, DP-SGD with uniform data sampling without replacement shares the same privacy guarantee as DP-SGD with Poisson subsampling.
III-C Zeroth-Order Optimization